# BitDecoding: Unlocking Tensor Cores for Long-Context LLMs with Low-Bit KV Cache

### v2

@@ Abstract
The rise of long-context Large Language Models (LLMs) amplifies memory and bandwidth demands during autoregressive decoding, as the Key–Value (KV) cache grows with each generated token. Low-bit KV-cache quantization (e.g., 4-bit or 2-bit) can reduce memory footprint while preserving accuracy, but existing systems suffer from slow decoding due to their exclusive reliance on CUDA cores, neglecting Tensor Cores—the primary source of compute on modern GPUs.
We present BitDecoding, a new long-context LLMs inference system with low-bit KV cache. BitDecoding enables efficient low-bit KV cache decoding by cooperatively leveraging CUDA Cores and Tensor Cores. It introduces methods for automatically inducing optimized layouts to exploit Tensor Cores, along with novel warp-level parallelization strategies for dequantization. For unified system support, BitDecoding includes a query transformation module supporting diverse attention variants, a quantization kernel to support both tensor-wise and channel-wise scaling used in various quantization algorithms with high performance, and a dequantization kernel with a software-defined pipeline to coordinate CUDA and Tensor Cores execution for mix-precision operations.
@@ I Introduction
The ability of Large Language Models (LLMs) to process long contexts [21, 6, 28] has unlocked new capabilities, such as book summarization [3], multi-modal understanding [33], and test-time scaling [10, 20]. However, these advancements come with significant memory and computational challenges, primarily due to the large size of the Key-Value (KV) cache in long-context scenarios. During autoregressive decoding, LLMs must repeatedly access this growing cache for each generated token, which increases memory usage and slows down decoding. The problem worsens with larger batch sizes, as the KV cache scales linearly with the number of concurrent queries. For example, a 7B model requires approximately 14GB GPU memory for its parameters, but with a 32K context length and a batch size of 8, the KV cache alone consumes 128GB GPU memory [11], creating a significant memory bottleneck.
@@
To address this growing bottleneck, KV cache quantization has emerged as a promising solution. By reducing the bit-width of the KV cache, quantization lowers memory overhead and improves overall efficiency. Recent quantization algorithms have shown that low-bit KV cache can retain high accuracy. QServe [15] demonstrates 4-bit KV cache improves throughput on models like LLaMA-3 and Qwen-1.5 while maintaining strong accuracy, even together with 4-bit weight and 8-bit activation. Further research [17, 12, 25] shows that 2-bit KV cache can achieve near fp16 accuracy. Kivi [17], for instance, incurs only a 0.6% accuracy drop on LongBench [2] with a 2-bit KV cache on LLaMA-2-7B-Chat. Recent studies [34, 27] explore 1-bit quantization for KV cache, maintaining acceptable accuracy under specific conditions. These results confirm that KV cache quantization strikes an effective balance between efficiency and accuracy, making it viable for long-context LLM deployment.
@@
Despite the memory savings, current system support for low-bit KV cache struggles to deliver the expected speedup. Previous implementations [17, 35, 15] remain preliminary and case-specific, with significant room for further systematic optimization. A major bottleneck lies in the overhead introduced by quantization and dequantization. Although the KV cache is low-bit, the query (Q) values and attention scores remain in high precision. This results in mixed-precision matrix multiplications (mpGEMM), which existing hardware does not natively support, requiring dequantization before multiplication. Previous mpGEMM kernels like Ladder [31] and Marlin [8] are designed for low-bit weights but cannot be directly applied to low-bit KV caches. This is because weights are static and stored offline, while KV caches are dynamic and generated online. In autoregressive decoding, each newly generated token requires quantization, packing, and dequantization of the low-bit KV cache, introducing significant overhead and complexity in GPU kernel design, as illustrated in Fig. 1.
@@
To address this, our insight is to leverage Tensor Cores for intensive matrix multiplications while efficiently utilizing CUDA cores for KV cache dequantization. Previous work either implemented with separated kernels or fused attention operations relied solely on CUDA cores, leaving Tensor Cores underutilized, as shown in Fig. 2. Our approach is based on three key observations: First, modern language models employ Grouped-Query Attention (GQA) and Multi-Query Attention (MQA), which share a group of keys across multiple queries, enabling Tensor Cores to accelerate dot products in the self-attention mechanism. Second, leveraging Tensor Cores can alleviate computational pressure on CUDA cores, enabling more efficient execution of dequantization. Finally, newer GPU architectures like Hopper support asynchronous execution on Tensor Cores [18], allowing the extra operations introduced by the low-bit KV cache to overlap with computation, further enhancing efficiency.
@@
To address the above challenges, we have designed and implemented BitDecoding, a high-performance long-context LLMs inference system with low-bit KV cache. The design of BitDecoding delivers several contributions essential for exploiting Tensor Cores, including: (i) inducing low-bit optimized layouts based on hardware instructions, (ii) aligning warps with residual buffer to saturate Tensor Cores, (iii) re-mapping layouts for faster dequantization, (iv) coordinating kernels for quantization and dequantization, and (v) unlocking warpgroup acceleration capabilities through the smart use of PTX-level instructions. In addition, we contribute new strategies for parallelizing GPU warps to mitigate dequantization overhead, including (i) efficient warp parallelism layout, and (ii) enhancing attention algorithms for fast warp synchronization leveraging the GPU memory hierarchy.
We further contribute implementation techniques in BitDecoding for LLMs inference, including: (i) a query transformation approach that enables efficient execution of diverse attention variants, allowing BitDecoding to be easily adopted in existing LLMs; (ii) a high-performance quantization kernel that supports both channel-wise and tensor-wise scaling, ensuring generality across quantization algorithms; and (iii) a dequantization kernel with a software-defined pipeline that coordinates CUDA and Tensor Cores for GEMM and dequantization, while overlapping data movement, including extra low-bit metadata.
BitDecoding is evaluated both at the kernel and end-to-end levels. At the kernel level, it outperforms FP16 FlashDecoding-v2 by up to $7.5\times$ (RTX 4090), $4.8\times$ (A100), and $8.9\times$ (H100), and surpassing Qserve by up to $4.3\times$. At the end-to-end model level, BitDecoding reduces single-batch decoding latency by $3\times$ on LLaMA-3.1-8B with a 128K sequence length and achieves over $4\times$ higher serving throughput than Qserve.
@@ II Background and Motivation
LLMs inference and low-bit KV cache. LLMs inference comprises two stages: (i) Prefill, which processes the prompt and computes Key (K) and Value (V) tensors for caching; and (ii) Decode, which updates the KV cache token-by-token for autoregressive generation. For a model with $n$ layers, $h_{kv}$ KV heads, and hidden size $d$, the KV cache requires $2\cdot 16\cdot n\cdot h_{kv}\cdot d\cdot b\cdot l$ bits (assuming FP16), where $b$ is the batch size and $l$ is the sequence length. Because this requirement grows linearly with both $b$ and $l$, the KV cache often dominates memory usage, especially for long-context and large-batch workloads. In batched inference, each sequence has an independent past context, so there is little batch-level parallelism or reuse when loading cached Keys and Values; consequently, KV-cache access is typically bound by memory bandwidth. These constraints have spurred extensive research and industrial efforts on lower-bit KV caches [17, 11, 34] to reduce memory footprint and improve throughput while preserving accuracy close to non-quantized baselines.
@@
Tensor Cores and CUDA cores on modern GPUs. When optimizing LLM inference and low-bit KV caches on GPUs, it is crucial to exploit both Tensor Cores and CUDA cores. Tensor Cores deliver the majority of compute FLOPS in modern GPUs but are specialized for matrix operations (e.g., GEMM), whereas CUDA cores provide more flexible vector, scalar, and control-flow capabilities at substantially lower peak FLOPS. For example, on the A100, Tensor Cores deliver up to 312 TFLOPS in FP16/BF16—far exceeding the 19.5 TFLOPS FP32 offered by CUDA cores. This FLOPS gap is even larger on newer GPUs such as Hopper and Blackwell.
@@
For fast LLM inference, substantial effort has gone into optimizing attention variants to exploit Tensor Cores. SOTA LLMs [16, 9, 32] increasingly adopt MQA [24] and GQA [1], which reduce memory bandwidth by reusing KV heads across multiple queries. This reuse increases arithmetic intensity and improves compute efficiency [26], aligning well with the high-throughput, matrix-centric design of Tensor Cores. Consequently, leveraging Tensor Cores is becoming essential for efficient inference in long-context and grouped-attention LLMs.
@@
Limitations of existing low-bit KV cache systems. To support low-bit KV caches for long-context LLM inference, a number of systems have been proposed [17, 15, 35]. However, they often leave GPUs underutilized, leading to suboptimal performance. We summarize the key reasons below.
@@
  • Attention with separated low-bit KV-cache kernels: The most straightforward approach, exemplified by Kivi [17], decomposes mixed-precision attention into multiple standalone kernels and embeds them in a non-fused attention implementation. This design is highly flexible and readily supports many attention variants [1, 24]. Yet the isolated launches repeatedly load and store intermediate data, inflate global-memory traffic, and break on-chip data reuse. The result is high launch overhead, increased memory bandwidth pressure, and lower effective throughput.
@@
  • Fused attention with low-bit KV-cache kernels on CUDA cores solely: Given the generality of CUDA cores for mixed-precision operations, a natural extension of FlashAttention-style fusion [5] is a CUDA-cores–only implementation of low-bit KV caches. While this outperforms non-fused designs, it still underutilizes Tensor Cores. In these systems, both dequantization and matrix operations (GEMV/GEMM) are executed on CUDA cores via fused multiply–add (FMA) instructions. Under mixed precision, CUDA cores must handle expensive dequantization (e.g., int4/8 $\rightarrow$ FP16/BF16), scaling, and element-wise ops—tasks that are memory-bound and consume instruction slots, register bandwidth, and L1/L2 capacity. This reduces occupancy and limits tile sizes, leaving fewer resources for the compute-heavy matrix multiplications. Consequently, running both dequantization and matmul on CUDA cores introduces significant overhead, especially for attention variants with higher arithmetic intensity.
@@ III-A Solution: Cooperative use of Tensor Cores & CUDA Cores
We expect significant benefits from this proposed solution. For example, the Hopper architecture introduces SM90 Tensor Core instructions and warp-specialized pipelines, yielding up to $6\times$ decoding speedups over prior implementations, whereas relying solely on SM80-generation instructions can reduce peak throughput by 35% [23]. Extending low-bit KV-cache decoding to leverage such hardware-accelerated instructions—e.g., warp-group–level scheduling and asynchronous execution—can approach order-of-magnitude gains (up to $10\times$) in decoding throughput. We anticipate even larger improvements on the upcoming Blackwell architecture, given the substantial increases in Tensor Cores capacity across GPU generations.
@@ III-B Open challenges
Challenge 1: Tensor Cores often suffer from low-bit layout mismatches. Aligning low-bit data layouts with Tensor Core requirements is difficult, especially in autoregressive generation where KV caches expand dynamically.
@@
Second, low-precision bitwidths exacerbate alignment issues. Although Tensor Core instructions require specific compute types, their rigid, interleaved register layout makes lower-precision data hard to match directly. Without a layout transform, the low-bit register layout becomes an invalid layout for MMA execution due to misalignment with the interleaved access patterns. As shown in Fig. 3(b), two FP16 values originally computed by Thread 0 (T0) may be quantized and packed as eight consecutive low-bit values in the KV cache; after dequantization, they no longer align with the expected Tensor Core register layout, yielding incorrect values.
@@
For example, under FlashAttention’s original warp partitioning, the additional dequantization (DQ) can substantially reduce throughput and Tensor Core utilization. As shown in Fig.4(a), FlashAttention assigns a single warp along the $N$ dimension to perform register-level softmax and the matrix multiplication $PV$, with $P$ stored in registers aligned to the Tensor Core layout. When DQ is inserted before the matmul, this strategy becomes inefficient: small warp tiles of $K$ or $V$ must traverse $N$ sequentially, so DQ frequently stalls the warp.
Nsight Compute profiling[19] in Fig.4(b) confirms that the added DQ overhead increases memory-access stalls and depresses compute throughput and Tensor Core utilization, consistent with prior observations[7].
Challenge 3: Lack of generalizable system optimizations for different low-bit KV-cache methods. Popular KV-cache quantization methods use diverse scaling granularities for the Key tensor—tensor-wise [35, 11] and channel-wise [17, 12]—which complicates building a unified system that supports them all. Online quantization and packing require reductions and element-wise transforms, adding nontrivial runtime overhead. Moreover, auxiliary metadata (scale and zero-point) increases memory traffic and, without careful scheduling, disrupts the load–compute pipeline. Prior mixed-precision kernel optimizations [31, 8] target static weight quantization and do not generalize to the dynamic, step-by-step nature of KV caches. To date, generalizable system-level optimization techniques for high-performance, low-bit KV-cache quantization are lacking.

### v3

@@ Abstract
The growth of long-context Large Language Models (LLMs) significantly increases memory and bandwidth pressure during autoregressive decoding due to the expanding Key–Value (KV) cache. While accuracy-preserving KV-cache quantization (e.g., 4-bit or 2-bit) reduces memory footprint, existing systems decode inefficiently by relying solely on CUDA cores, underutilizing Tensor Cores—the dominant compute resource on GPUs.
We present BitDecoding, the first inference system to efficiently decode low-bit KV caches by cooperatively leveraging CUDA cores and Tensor Cores. BitDecoding smartly induces Tensor-Core-friendly layouts, introduces warp-level dequantization parallelism, and provides unified system support through query transformation, high-performance tensor- and channel-wise quantization, and a software-pipelined dequantization kernel enabling mixed-precision execution. Architecture-aware optimizations further leverage Hopper’s warpgroup tensor instructions and Blackwell’s NVFP4 (MXFP4) tensor formats.
@@ I Introduction
The ability of Large Language Models (LLMs) to process long contexts [yarn, longrope, gemini] has unlocked new capabilities, such as book summarization [booookscore], multi-modal understanding [mm], and test-time scaling [deepseek-r1, openai_o3_mini]. However, these advancements come with significant memory and computational challenges, primarily due to the large size of the Key-Value (KV) cache in long-context scenarios. During autoregressive decoding, LLMs must repeatedly access this growing cache for each generated token, which increases memory usage and slows down decoding. The problem worsens with larger batch sizes, as the KV cache scales linearly with the number of concurrent queries. For example, a 7B model requires approximately 14GB GPU memory for its parameters, but with a 32K context length and a batch size of 8, the KV cache alone consumes 128GB GPU memory [kvquant], creating a significant memory bottleneck.
@@
To address this growing bottleneck, KV cache quantization has emerged as a promising solution. By reducing the bit-width of the KV cache, quantization lowers memory overhead and improves overall efficiency. Recent quantization algorithms have shown that low-bit KV cache can retain high accuracy. QServe [qserve] demonstrates 4-bit KV cache improves throughput on models like LLaMA-3 and Qwen-1.5 while maintaining strong accuracy, even together with 4-bit weight and 8-bit activation. Further research [kivi, gear, rotatekv] shows that 2-bit KV cache can achieve near fp16 accuracy. Kivi [kivi], for instance, incurs only a 0.6% accuracy drop on LongBench [bai2024longbench] with a 2-bit KV cache on LLaMA-2-7B-Chat. Recent studies [1bitkv, asymkv] explore 1-bit quantization for KV cache, maintaining acceptable accuracy under specific conditions. These results confirm that KV cache quantization strikes an effective balance between efficiency and accuracy, making it viable for long-context LLM deployment.
@@
Despite the memory savings, current system support for low-bit KV cache struggles to deliver the expected speedup. Previous implementations [kivi, atom, qserve] remain preliminary and case-specific, with significant room for further systematic optimization. A major bottleneck lies in the overhead introduced by quantization and dequantization. Although the KV cache is low-bit, the query (Q) values and attention scores remain in high precision. This results in mixed-precision matrix multiplications (mpGEMM), which existing hardware does not natively support, requiring dequantization before multiplication. Previous mpGEMM kernels like Ladder [ladder] and Marlin [marlin] are designed for low-bit weights but cannot be directly applied to low-bit KV caches. This is because weights are static and stored offline, while KV caches are dynamic and generated online. In autoregressive decoding, each newly generated token requires quantization, packing, and dequantization of the low-bit KV cache, introducing significant overhead and complexity in GPU kernel design, as illustrated in Fig. 1.
@@
To address this, our insight is to leverage Tensor Cores for intensive matrix multiplications while efficiently utilizing CUDA cores for KV cache dequantization. Previous work either implemented with separated kernels or fused attention operations relied solely on CUDA cores, leaving Tensor Cores underutilized, as shown in Fig. 2. Our approach is based on three key observations: First, modern language models employ Grouped-Query Attention (GQA) and Multi-Query Attention (MQA), which share a group of keys across multiple queries, enabling Tensor Cores to accelerate dot products in the self-attention mechanism. Second, leveraging Tensor Cores can alleviate computational pressure on CUDA cores, enabling more efficient execution of low-bit operations. Finally, newer GPU architectures provide distinct mechanisms: Hopper’s support for asynchronous execution and warp specialization allows low-bit operations to overlap with computation [luo2024benchmarking], while Blackwell’s native support for low-precision formats (e.g., MXFP4) reduces these overheads by minimizing the need for on-the-fly data conversion.
@@
To address the above challenges, we have designed and implemented BitDecoding, a high-performance long-context LLMs inference system with low-bit KV cache. The design of BitDecoding delivers several contributions essential for exploiting Tensor Cores, including: (i) inducing low-bit optimized layouts based on hardware instructions, (ii) aligning warps with residual buffer to saturate Tensor Cores, (iii) re-mapping layouts for faster dequantization, and (iv) coordinating kernels for quantization and dequantization. In addition, we contribute new strategies for parallelizing GPU warps to mitigate low-bit operations overhead, including (i) efficient warp parallelism layout, and (ii) enhancing attention algorithms for fast warp synchronization leveraging the GPU memory hierarchy.
We further contribute implementation techniques in BitDecoding for LLMs inference, including: (i) a query transformation approach that enables efficient execution of diverse attention variants, allowing BitDecoding to be easily adopted in existing LLMs; (ii) a high-performance quantization kernel that supports both channel-wise and tensor-wise scaling, ensuring generality across quantization algorithms; and (iii) a dequantization kernel with a software-defined pipeline that coordinates CUDA and Tensor Cores for GEMM and dequantization, while overlapping data movement, including extra low-bit metadata; furthermore, BitDecoding incorporates architecture-specific optimizations that unlock Hopper’s warpgroup tensor operations and Blackwell’s native low-precision tensor formats to maximize decoding performance on the latest GPU generations.
[…]
@@ II Background and Motivation
LLMs inference and low-bit KV cache. LLMs inference comprises two stages: (i) Prefill, which processes the prompt and computes Key (K) and Value (V) tensors for caching; and (ii) Decode, which updates the KV cache token-by-token for autoregressive generation. For a model with $n$ layers, $h_{kv}$ KV heads, and hidden size $d$, the KV cache requires $2\cdot 16\cdot n\cdot h_{kv}\cdot d\cdot b\cdot l$ bits (assuming FP16), where $b$ is the batch size and $l$ is the sequence length. Because this requirement grows linearly with both $b$ and $l$, the KV cache often dominates memory usage, especially for long-context and large-batch workloads. In batched inference, each sequence has an independent past context, so there is little batch-level parallelism or reuse when loading cached Keys and Values; consequently, KV-cache access is typically bound by memory bandwidth. These constraints have spurred extensive research and industrial efforts on lower-bit KV caches [kivi, kvquant, 1bitkv] to reduce memory footprint and improve throughput while preserving accuracy close to non-quantized baselines.
@@
Tensor Cores and CUDA cores on modern GPUs. When optimizing LLM inference and low-bit KV caches on GPUs, it is crucial to exploit both Tensor Cores and CUDA cores. Tensor Cores deliver the majority of compute FLOPS in modern GPUs but are specialized for matrix operations (e.g., GEMM), whereas CUDA cores provide more flexible vector, scalar, and control-flow capabilities at substantially lower peak FLOPS. For example, on the A100, Tensor Cores deliver up to 312 TFLOPS in FP16/BF16—far exceeding the 19.5 TFLOPS FP32 offered by CUDA cores.
This performance gap has widened significantly in recent generations. The Hopper architecture introduces Warpgroup Matrix Multiply-Accumulate (WGMMA) instructions and warp-specialized pipelines to maximize asynchronous execution efficiency. The Blackwell architecture further exacerbates this disparity by supporting native micro-scaling formats (e.g., MXFP4, NVFP4), delivering up to 20 PFLOPS.
@@
For fast LLM inference, substantial effort has gone into optimizing attention variants to exploit Tensor Cores. SOTA LLMs [liu2024deepseek, llama3, qwen3] increasingly adopt MQA [mqa] and GQA [gqa], which reduce memory bandwidth by reusing KV heads across multiple queries. This reuse increases arithmetic intensity and improves compute efficiency [gta], aligning well with the high-throughput, matrix-centric design of Tensor Cores. Consequently, leveraging Tensor Cores is becoming essential for efficient inference in long-context and grouped-attention LLMs.
@@
Limitations of existing low-bit KV cache systems. To support low-bit KV caches for long-context LLM inference, a number of systems have been proposed [kivi, qserve, atom]. However, they often leave GPUs underutilized, leading to suboptimal performance. We summarize the key reasons below.
@@
  • Attention with separated low-bit KV-cache kernels: The most straightforward approach, exemplified by Kivi [kivi], decomposes mixed-precision attention into multiple standalone kernels and embeds them in a non-fused attention implementation. This design is highly flexible and readily supports many attention variants [gqa, mqa]. Yet the isolated launches repeatedly load and store intermediate data, inflate global-memory traffic, and break on-chip data reuse. The result is high launch overhead, increased memory bandwidth pressure, and lower effective throughput.
@@
  • Fused attention with low-bit KV-cache kernels on CUDA cores solely: Given the generality of CUDA cores for mixed-precision operations, a natural extension of FlashAttention-style fusion [dao2023flashattention2] is a CUDA-cores–only implementation of low-bit KV caches. While this outperforms non-fused designs, it still underutilizes Tensor Cores. In these systems, both dequantization and matrix operations (GEMV/GEMM) are executed on CUDA cores via fused multiply–add (FMA) instructions. Under mixed precision, CUDA cores must handle expensive dequantization (e.g., int4/8 $\rightarrow$ FP16/BF16), scaling, and element-wise ops—tasks that are memory-bound and consume instruction slots, register bandwidth, and L1/L2 capacity. This reduces occupancy and limits tile sizes, leaving fewer resources for the compute-heavy matrix multiplications. Consequently, running both dequantization and matmul on CUDA cores introduces significant overhead, especially for attention variants with higher arithmetic intensity.
@@ III-A Solution: Cooperative use of Tensor Cores & CUDA Cores
We expect significant benefits from this proposed solution. For example, by enabling low-bit decoding that builds on FlashAttention-3 (FA-3) [fa3], we can leverage SM90-specific features—such as warp-specialized pipelines—that yield up to $6\times$ speedups over prior implementations, avoiding the 35% throughput penalty associated with legacy SM80 instructions. Furthermore, this design anticipates the architectural capabilities of Blackwell, where native support for low-precision formats will drive even more substantial throughput improvements.
@@ III-B Open challenges
Challenge 1: Tensor Cores often suffer from low-bit layout mismatches. Aligning low-bit data layouts with Tensor Cores requirements is difficult, especially in autoregressive generation where KV caches expand dynamically.
@@
Second, low-precision bitwidths exacerbate alignment issues. Although Tensor Cores instructions require specific compute types, their rigid, interleaved register layout makes lower-precision data hard to match directly. Without a layout transform, the low-bit register layout becomes an invalid layout for MMA execution due to misalignment with the interleaved access patterns. As shown in Fig. 3(b), two FP16 values originally computed by Thread 0 (T0) may be quantized and packed as eight consecutive low-bit values in the KV cache; after unpacking and dequantization, they no longer align with the expected Tensor Core register layout, yielding incorrect values. Even with native low-precision formats in Blackwell, hardware support remains limited, especially for the KV cache, which still depends on continuous quantization and packing; software must therefore carefully handle low-precision values and micro-scaling factors [nvidia_triton_blackwell_2025].
@@
Under FlashAttention’s original warp partitioning, the additional dequantization (DQ) can substantially reduce throughput and Tensor Core utilization. As shown in Fig. 4(a), FlashAttention assigns a single warp along the $N$ dimension to perform register-level softmax and the matrix multiplication $PV$, with $P$ stored in registers aligned to the Tensor Core layout. When DQ is inserted before the matmul, this strategy becomes inefficient: small warp tiles of $K$ or $V$ must traverse $N$ sequentially, so DQ frequently stalls the warp. Nsight Compute profiling[NVIDIA_Nsight_Compute] in Fig. 4(b) confirms that the added DQ overhead increases memory-access stalls and depresses compute throughput and Tensor Cores utilization, consistent with prior observations[fan2025warpdrive].
Furthermore, native low-precision formats introduce their own overhead despite eliminating dequantization. Specifically, to utilize low-precision Tensor Cores for the second matrix multiplication ($PV$), the probability matrix $P$ must be dynamically re-quantized after the softmax operation: $P_{f16}=\mathrm{softmax}(Q_{f4}K_{f4}^{T}),\quad O_{f16}=\mathbf{Quant}(P_{f16})V_{f4}$. This on-the-fly quantization creates a new computational bottleneck that can similarly stall Tensor Cores execution.
Challenge 3: Lack of generalizable system optimizations for different low-bit KV-cache methods. Popular KV-cache quantization methods use diverse scaling granularities for the Key tensor—tensor-wise [atom, kvquant] and channel-wise [kivi, gear]—which complicates building a unified system that supports them all. Online quantization and packing require reductions and element-wise transforms, adding nontrivial runtime overhead. Moreover, auxiliary metadata (scale and zero-point) increases memory traffic and, without careful scheduling, disrupts the load–compute pipeline. Prior mixed-precision kernel optimizations [ladder, marlin] target static weight quantization and do not generalize to the dynamic, step-by-step nature of KV caches. To date, generalizable system-level optimization techniques for high-performance, low-bit KV-cache quantization are lacking.